# Detection of Correlated Random Vectors

### v1

@@ 5.1 Lower bounds
We follow the steps of the proofs from the one dimensional case. Note that (23) and (25) are true regardless to the structure of our problem and therefore holds for generalized model as well. We take on the same approach for evaluating the second moment of the likelihood ratio by decomposing it orthogonal components and using Parseval’s equality. We consider the Hilbert space $L^{2}({\cal H}_{0})$ of all random variables over $\mathbb{R}^{n\times}\times\mathbb{R}^{n\times d}$ with the inner product defined in (26). As in the one-dimensional case, the set of multivariate Hermite polynomials $\left\{H_{\alpha,\beta}\right\}_{\alpha,\beta}$ compose an orthonormal basis, only that this time $\alpha,\beta\in\mathbb{N}^{n\times d}$ are $n\times d$ coefficient matrices (rather than vectors), and
@@
For an $n\times d$ integer matrix $\alpha$ define $|\alpha|=\sum_{i,j}\alpha_{i,j}$ to be the $L^{1}$ norm and $\mathsf{RS}(\alpha)$ to be the row support of $\alpha$, which is, the set of indices of the non-zero rows of $\alpha$:
@@
We also generalize the definition of the equivalence relation presented in Section 4 to the integer matrices. We say that $\alpha\equiv\beta$ if there exists a permutation $\sigma\in{\mathbb{S}}_{n}$ such that $\beta$ is obtained by permuting the rows of $\alpha$ according to $\sigma$. We denote the equivalence class of $\alpha$ by $[\alpha]$.
As mentioned, we use Parseval’s equality and obtain
@@
$=\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}\alpha,\beta\in\mathbb{N}^{n\times d}\\ |\alpha|+|\beta|=m\end{subarray}}\left\langle H_{\alpha,\beta}(\mathsf{X},\mathsf{Y}),{\cal L}(\mathsf{X},\mathsf{Y})\right\rangle_{{\cal H}_{0}}^{2}.$ (125)
Our goal now is to give an exact expression the orthogonal projection coefficients $\left\langle H_{\alpha,\beta}(\mathsf{X},\mathsf{Y}),{\cal L}(\mathsf{X},\mathsf{Y})\right\rangle_{{\cal H}_{0}}$, which we calculate in a similar fashion as in Lemma 1.
@@
We start the proof by repeating the steps of the proof of Lemma 1, using the independence of the pairs $(\mathsf{X_{i,j}},\mathsf{Y}_{\sigma(i),j})_{i,j}$ given the permutation $\sigma$ and the set ${\cal K}$, we have
@@
From our model definition, conditioned on $\sigma$ and ${\cal K}$, for all $i\in[n]$, if $i\notin{\cal K}$ then $\mathsf{X}_{i,j}$ and $\mathsf{Y}_{\sigma(i),j}$ are independent. Thus, if $\alpha_{i,j}$ or $\beta_{\sigma(i),j}$ are non-zero
@@
where the last equality follows since distinct Hermite polynomials are orthogonal under $\left\langle\cdot,\cdot\right\rangle_{{\cal H}_{0}}$ and therefore
@@
if $\beta_{\sigma(i),j}\neq 0$. Thus, given fixed $\sigma$ and ${\cal K}$ the product
@@
can be non-zero only if $\alpha_{i,j}=\beta_{\sigma(i),j}=0$ for any $i\neq{\cal K}$ and $j\in[d]$.
On the other hand, for fixed $\sigma$, ${\cal K}$, $i\in{\cal K}$ and $j\in[d]$, by the proof of Lemma 1 we have
@@
For fixed $\alpha\equiv\beta$, note that ${\mathbb{P}}\left[\mathsf{RS}(\alpha)\subseteq{\cal K}\right]$ is non-zero only if $|\mathsf{RS}(\alpha)|\leq k$, in which case, since ${\cal K}$ is uniform over $\binom{[n]}{k}$,
@@
We carry on with our analysis of $\mathbb{E}_{{\cal H}_{0}}[{\cal L}(\mathsf{X},\mathsf{Y})^{2}]$ form (125), plugging in the orthogonal projection coefficients calcaulted in Lemma 3:
$\mathbb{E}_{{\cal H}_{0}}\left[{\cal L}(\mathsf{X},\mathsf{Y})^{2}\right]$ $=\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}\alpha,\beta\in\mathbb{N}^{n\times d}\\ |\alpha|+|\beta|=m\end{subarray}}\left\langle H_{\alpha,\beta}(\mathsf{X},\mathsf{Y}),{\cal L}(\mathsf{X},\mathsf{Y})\right\rangle_{{\cal H}_{0}}^{2}$ (152)
$=\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}\alpha,\beta\in\mathbb{N}^{n\times d}\\ |\alpha|+|\beta|=m\end{subarray}}\mathbbm{1}\left\{\alpha\equiv\beta,|\mathsf{RS}(\alpha)|\leq k\right\}\frac{\rho^{2|\alpha|}}{\left|[\alpha]\right|^{2}}\prod_{i=0}^{\left|\mathsf{RS}(\alpha)-1\right|}\left(\frac{k-i}{n-i}\right)^{2}$ (155)
$=\sum_{m=0}^{\infty}\sum_{\ell=0}^{k}\sum_{\begin{subarray}{c}\alpha\in\mathbb{N}^{n\times d}\\ |\alpha|=m\\ |\mathsf{RS}(\alpha)|=\ell\end{subarray}}\sum_{\beta\equiv\alpha}\frac{\rho^{2m}}{\left|[\alpha]\right|^{2}}\prod_{i=0}^{\ell-1}\left(\frac{k-i}{n-i}\right)^{2}$ (159)
$=\sum_{m=0}^{\infty}\rho^{2m}\sum_{\ell=0}^{k}\prod_{i=0}^{\ell-1}\left(\frac{k-i}{n-i}\right)^{2}\sum_{\begin{subarray}{c}\alpha\in\mathbb{N}^{n\times d}\\ |\alpha|=m\\ |\mathsf{RS}(\alpha)|=\ell\end{subarray}}\sum_{\beta\equiv\alpha}\frac{1}{\left|[\alpha]\right|^{2}}$ (163)
$=\sum_{m=0}^{\infty}\rho^{2m}\sum_{\ell=0}^{k}\prod_{i=0}^{\ell-1}\left(\frac{k-i}{n-i}\right)^{2}\sum_{\begin{subarray}{c}\alpha\in\mathbb{N}^{n\times d}\\ |\alpha|=m\\ |\mathsf{RS}(\alpha)|=\ell\end{subarray}}\frac{1}{\left|[\alpha]\right|}$ (167)
$=\sum_{m=0}^{\infty}\rho^{2m}\sum_{\ell=0}^{k}\prod_{i=0}^{\ell-1}\left(\frac{k-i}{n-i}\right)^{2}\left|\left\{[\alpha]~{}\bigg{|}\begin{subarray}{c}\alpha\in\mathbb{N}^{n\times d}\\ |\alpha|=m\\ |\mathsf{RS}(\alpha)|=\ell\end{subarray}\right\}\right|.$ (171)
@@
We now turn to estimate the number of equivalence classes $\mathsf{[}\alpha]$, for $\alpha\in\mathbb{N}^{n\times d}$, $|\alpha|=m$ and $|\mathsf{RS}(\alpha)|=\ell$ for fixed $\ell\leq k$. To that end, as in the one dimensional case, we point out a correspondence between equivalence classes and $d$-dimensional integer distribution functions with certain properties. For $\alpha\in\mathbb{N}^{n\times d}$, consider the function $p_{\alpha}:\mathbb{N}^{d}\to\mathbb{N}$ such that $p_{\alpha}(v)$ is the number of rows of $\alpha$ which are equal to $v\in\mathbb{N}^{d}$. We also note that $|\alpha|=m$ and $|\mathsf{RS}(\alpha)|=\ell$ if and only if
@@
$\leq\sum_{p_{1}\in\mathsf{Par}(m,\ell)}\prod_{q\in\mathbb{N}}\binom{q+d-1}{d-1}^{p_{1}(q)}$ (174)
$\leq d^{m}|\mathsf{Par}(m,\ell)|,$ (175)
where $p_{1}$ is a partition of the integer $m$ to exactly $\ell$ parts, which we think of as one dimensional integer distribution function ($p_{1}(q)$ is the number of parts of size $q$ in a given partition).
@@
We note that any $d$-dimensional distribution function $p$ which satisfies (172) corresponds to an unordered list of $\ell$ vectors in $\mathbb{N}^{d}$, such that the sum over all entries of all vectors together is $m$. For a given vector $v\in\mathbb{N}^{d}$, $p(v)$ represents the number of occurrences of the vector $v$ in that list. We upper bound the number of such lists
$\sum_{p_{1}\in\mathsf{Par}(m,\ell)}\prod_{q\in\mathbb{N}}\binom{q+d-1}{d-1}^{p_{1}(q)},$ (176)
by considering the following procedure of generating all possible lists:
We first go over all possible norms of vectors $v$ in that list. Since the summation of these weights is exactly $m$, we should go over all partitions of the integer $m$ to exactly $\ell$ elements. Having the weights of the vectors in the list fixed, we upper bound the number of ways to choose vectors with these exact weights. For a partition part with weight $q$, the number of vectors in $\mathbb{N}^{d}$ with weight $q$ is exactly as the number of ways to distribute $q$ elements into $d$ cells, which is known to be $\binom{q+d-1}{d-1}$. For a fixed partition $p_{1}$, there are $p_{1}(q)$ parts with weight $q$, and therefore, there are at most $\binom{q+d-1}{d-1}^{p_{1}(q)}$ ways to generate $d$-dimensional integer vectors with such norm. Taking into account all of the distinct parts of the partition $p_{1}$ we upper bound the number ways to generate appropriate $d$-dimensional integer vectors we by
@@
For the second upper bound, we note that the $\binom{q+d-1}{d-1}$, the number of ways to distribute $q$ elements into $d$ cells, can always be upper bounded by $d^{q}$. When considering each of the $q$ elements has distinct item, it has $d$ optional cells, and therefore, together, the $q$ elements has $d^{q}$ optional distributions along $d$ cells. We therefore have
@@
For the lower bound, we note that for a fixed $q$, the expression in (177) may over count the number of possible ways to choose $p_{1}(q)$ unordered $d$-dimension vectors whose norms are $q$, since different orderings of such vectors are counted multiple times. Thus, for we lower bound the number possible ways to choose $p_{1}(q)$ unordered $d$-dimension vectors whose norms are $q$ by dividing in $p_{1}(q)!$ which is maximal number of multiple counts for each set of $p_{1}(q)$ vectors. The rest of the analysis is carried out similarly to the upper bound. ∎
@@
Proof of Theorem 6 and Theorem 5.
We continue from (171), and use Lemma 4:
$\mathbb{E}_{{\cal H}_{0}}\left[{\cal L}(\mathsf{X},\mathsf{Y})^{2}\right]$ $=\sum_{m=0}^{\infty}\rho^{2m}\sum_{\ell=0}^{k}\prod_{i=0}^{\ell-1}\left(\frac{k-i}{n-i}\right)^{2}\left|\left\{[\alpha]~{}\bigg{|}\begin{subarray}{c}\alpha\in\mathbb{N}^{n\times d}\\ |\alpha|=m\\ |\mathsf{RS}(\alpha)|=\ell\end{subarray}\right\}\right|$ (186)
$=\sum_{m=0}^{\infty}\rho^{2m}\sum_{\ell=0}^{k}\prod_{i=0}^{\ell-1}\left(\frac{k-i}{n-i}\right)^{2}\left|\left\{[\alpha]~{}\bigg{|}\begin{subarray}{c}\alpha\in\mathbb{N}^{n\times d}\\ |\alpha|=m\\ |\mathsf{RS}(\alpha)|=\ell\end{subarray}\right\}\right|$ (190)
$=\sum_{m=0}^{\infty}\rho^{2m}\sum_{\ell=0}^{k}\prod_{i=0}^{\ell-1}\left(\frac{k-i}{n-i}\right)^{2}d^{m}\left|\mathsf{Par}(m,\ell)\right|$ (191)
$\overset{(a)}{\leq}\sum_{m=0}^{\infty}\left(d\rho^{2}\right)^{m}\sum_{\ell=0}^{k}\left(\frac{k}{n}\right)^{2\ell}\left|\mathsf{Par}(m,\ell)\right|$ (192)
$\overset{(b)}{\leq}\sum_{m=0}^{\infty}\left(d\rho^{2}\right)^{m}\sum_{\ell=0}^{k}\left(\frac{k}{n}\right)^{2}\left|\mathsf{Par}(m,\ell)\right|$ (193)
$=\left(\frac{k}{n}\right)^{2}\sum_{m=0}^{\infty}\left(d\rho^{2}\right)^{m}\sum_{\ell=0}^{k}\left|\mathsf{Par}(m,\ell)\right|$ (194)
$=\left(\frac{k}{n}\right)^{2}\sum_{m=0}^{\infty}\left(d\rho^{2}\right)^{m}\left|\mathsf{Par}(m,\leq_{k})\right|,$ (195)
where $(a)$ follows since the function $\frac{k-i}{n-i}$ is monotone decreasing in $i$ as long as $n>k$, and $(b)$ follows since $|\mathsf{Par(m,0)}|=0$ for any positive integer and $(k/n)^{2\ell}$ is decreasing with $\ell$.
We recall the well-known generating function of integer partitions to at most $k$ elements (see, for example, [FS09, Chapter 1.3])
@@
where the infinite sum converges for $|z|<1$. Thus, for any $\rho^{2}<\frac{1}{d}$ the infinite sum in (195) converge and we have
$\mathbb{E}_{{\cal H}_{0}}[{\cal L}(\mathsf{X},\mathsf{Y})]$ $\leq\left(\frac{k}{n}\right)^{2}\sum_{m=0}^{\infty}|\mathsf{Par}(m,\leq_{k})|(d\rho^{2})^{m}$ (197)
$=\left(\frac{k}{n}\right)^{2}\prod_{i=1}^{k}\frac{1}{1-(d\rho^{2})^{i}}.$ (198)
We conclude the repeating the steps from the proofs of Theorem 1 and Theorem 2, using Cauchy-Schwartz inequality (25) for impossibility of weak detection and Lemma 4 for the impossibility of strong detection. ∎
Let us now compare our proof technique with the one of [KN22b, EH23, PH23], where a straightforward analysis is carried out. Repeating their steps and generalizing for the partly correlated case, one can show that $\mathbb{E}_{{\cal H}_{0}}[{\cal L}(\mathsf{X},\mathsf{Y})^{2}]$ can similarly be expressed using cycle enumerators of random permutations.
@@
The second moment of the likelihood ratio for the detection problem defined in (100) is given by
@@
In the case where $k=n$ (fully correlated databases), where the model reduces to the one studied in [EH23, KN22b], $\mathsf{N}_{\ell}(\sigma,{\cal K})$ becomes $N_{\ell}(\sigma)$, the number of cycles of length $\ell$ of a random permutation. The analysis of (199) is already involved and requires delicate case by case analysis for the different regimes (of $n,d$ and $\rho$), using a Poisson approximation of the random variables $(N_{\ell}(\sigma))_{\ell}$. Using our method of calculating $\mathbb{E}_{{\cal H}_{0}}[{\cal L}(\mathsf{X},\mathsf{Y})^{2}]$ by polynomial decomposition on the other hand, the results of [EH23] for the cases where the $d=\Omega(1)$ are reproved, for all regimes at once.
In the case where $k<n$ (partial correlation), it becomes unclear how to evaluate the expression in (199). Conditioning on the size of the intersection ${\cal K}\cap[k]$ (which we denote by $i$), it is possible to express (199) as a weighted average of elements of the form
@@
where $\mathsf{N}_{\ell}(\sigma,[i])$ is the size of the number of cycles of $\sigma$ intersecting the interval $[i]$ in exactly $\ell$ points. It seems likely that a Poisson approximation for the variables $(\mathsf{N}_{\ell}(\sigma,[i]))_{\ell,i}$ exists and may be used in order to estimate $\mathbb{E}_{{\cal H}_{0}}[{\cal L}(\mathsf{X},\mathsf{Y})^{2}]$. However, proving such Poisson approximation theorems is highly non-trivial, and has its own intellectual merit. We therefore leave it for future research.

### v3

@@ 5.1 Lower bounds
As in the proof for the $d=1$ case, throughout this entire section, with some abuse of notation, we use lower-case Greek letters (such as, $\alpha$) for integer matrices, and the dimensions of these matrices are mentioned of understood by context. The prove strategy of Theorems 56 is the same as that of Theorems 12. We thereby repeat the exact same steps described in the proof outline presented in Section 4.2. However, we would like to emphasize that despite this similarity, a more refined analysis is required for the more complicated generalized multidimensional partly-correlated scenario. Note that (29) and (32) are true regardless of the structure of our problem and therefore hold for the generalized model as well. Accordingly, we follow the same approach of evaluating the second moment of the likelihood ratio by representing it as a linear combination of basis functions, and then use Parseval’s identity. We consider the Hilbert space $L^{2}({\cal H}_{0})$ of all random variables over $\mathbb{R}^{n\times d}\times\mathbb{R}^{n\times d}$, with the inner product defined in (16). As in the one-dimensional case, the set of multivariate Hermite polynomials $\left\{H_{\alpha,\beta}\right\}_{\alpha,\beta}$ form an orthonormal basis, but now $\alpha,\beta\in\mathbb{N}^{n\times d}$ are $n\times d$ coefficient matrices (rather than vectors), and
@@
For an $n\times d$ integer matrix $\alpha$ define $|\alpha|=\sum_{i,j}\alpha_{i,j}$ to be its $L^{1}$ norm, and $\mathsf{RS}(\alpha)$ to be the row support of $\alpha$, that is, the set of indices of the non-zero rows of $\alpha$,
@@
We also generalize the definition of the equivalence relation we defined in Section 4 for integer vectors to integer matrices. We say that $\alpha\equiv\beta$ if there exists a permutation $\sigma\in{\mathbb{S}}_{n}$ such that $\beta$ is obtained by permuting the rows of $\alpha$ according to $\sigma$. We denote the equivalence class of $\alpha$ by $[\alpha]$. Now, using Parseval’s identity, we obtain,
@@
$=\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}(\alpha,\beta)\in(\mathbb{N}^{n\times d})^{2}\\ |\alpha|+|\beta|=m\end{subarray}}\left\langle H_{\alpha,\beta}(\mathsf{X},\mathsf{Y}),{\cal L}(\mathsf{X},\mathsf{Y})\right\rangle_{{\cal H}_{0}}^{2}.$ (117)
Our goal now is to give an exact expression for the orthogonal projection coefficients $\left\langle H_{\alpha,\beta}(\mathsf{X},\mathsf{Y}),{\cal L}(\mathsf{X},\mathsf{Y})\right\rangle_{{\cal H}_{0}}$. This is done in a similar fashion to Lemma 1.
@@
We start the proof by repeating the steps of the proof of Lemma 1. Given $\sigma$ and ${\cal K}$, the pairs $(\mathsf{X_{i,j}},\mathsf{Y}_{\sigma(i),j})_{i,j}$ are statistically independent random variables, and thus,
@@
Now, recall that conditioned on $\sigma$ and ${\cal K}$, for all $i\in[n]$, if $i\notin{\cal K}$, then $\mathsf{X}_{i,j}$ and $\mathsf{Y}_{\sigma(i),j}$ are independent. Thus, if $\alpha_{i,j}$ or $\beta_{\sigma(i),j}$ are non-zero, we get
@@
where the last equality follows since distinct Hermite polynomials are orthogonal under $\left\langle\cdot,\cdot\right\rangle_{{\cal H}_{0}}$, and therefore
@@
if $\beta_{\sigma(i),j}\neq 0$. Thus, on one hand, given $\sigma$ and ${\cal K}$, we obtain that the product
@@
can be non-zero only if $\alpha_{i,j}=\beta_{\sigma(i),j}=0$ for any $i\neq{\cal K}$ and $j\in[d]$. On the other hand, given $\sigma$, ${\cal K}$, $i\in{\cal K}$ and $j\in[d]$, by the proof of Lemma 1, we have
@@
For a fixed $\alpha\equiv\beta$, note that ${\mathbb{P}}\left[\mathsf{RS}(\alpha)\subseteq{\cal K}\right]$ is non-zero only if $|\mathsf{RS}(\alpha)|\leq k$, in which case, since ${\cal K}$ is uniform over $\binom{[n]}{k}$,
@@
We carry on with our analysis of $\mathbb{E}_{{\cal H}_{0}}[{\cal L}(\mathsf{X},\mathsf{Y})^{2}]$ in (117). Plugging in the orthogonal projection coefficients calculated in Lemma 3, we obtain,
$\mathbb{E}_{{\cal H}_{0}}\left[{\cal L}(\mathsf{X},\mathsf{Y})^{2}\right]$ $=\sum_{m=0}^{\infty}\sum_{\begin{subarray}{c}(\alpha,\beta)\in(\mathbb{N}^{n\times d})^{2}\\ |\alpha|+|\beta|=m\end{subarray}}\left\langle H_{\alpha,\beta}(\mathsf{X},\mathsf{Y}),{\cal L}(\mathsf{X},\mathsf{Y})\right\rangle_{{\cal H}_{0}}^{2}$ (138)
$=1+\sum_{m=1}^{\infty}\sum_{\begin{subarray}{c}(\alpha,\beta)\in(\mathbb{N}^{n\times d})^{2}\\ |\alpha|+|\beta|=m\end{subarray}}\left\langle H_{\alpha,\beta}(\mathsf{X},\mathsf{Y}),{\cal L}(\mathsf{X},\mathsf{Y})\right\rangle_{{\cal H}_{0}}^{2}$ (139)
$=1+\sum_{m=1}^{\infty}\sum_{\begin{subarray}{c}(\alpha,\beta)\in(\mathbb{N}^{n\times d})^{2}\\ |\alpha|+|\beta|=m\end{subarray}}\mathbbm{1}\left\{\alpha\equiv\beta,|\mathsf{RS}(\alpha)|\leq k\right\}\frac{\rho^{2|\alpha|}}{\left|[\alpha]\right|^{2}}\prod_{i=0}^{\left|\mathsf{RS}(\alpha)-1\right|}\left(\frac{k-i}{n-i}\right)^{2}$ (140)
$=1+\sum_{m=1}^{\infty}\sum_{\ell=0}^{k}\sum_{\begin{subarray}{c}\alpha\in\mathbb{N}^{n\times d}\\ |\alpha|=m\\ |\mathsf{RS}(\alpha)|=\ell\end{subarray}}\sum_{\beta\equiv\alpha}\frac{\rho^{2m}}{\left|[\alpha]\right|^{2}}\prod_{i=0}^{\ell-1}\left(\frac{k-i}{n-i}\right)^{2}$ (141)
$=1+\sum_{m=1}^{\infty}\rho^{2m}\sum_{\ell=0}^{k}\prod_{i=0}^{\ell-1}\left(\frac{k-i}{n-i}\right)^{2}\sum_{\begin{subarray}{c}\alpha\in\mathbb{N}^{n\times d}\\ |\alpha|=m\\ |\mathsf{RS}(\alpha)|=\ell\end{subarray}}\sum_{\beta\equiv\alpha}\frac{1}{\left|[\alpha]\right|^{2}}$ (142)
$=1+\sum_{m=1}^{\infty}\rho^{2m}\sum_{\ell=0}^{k}\prod_{i=0}^{\ell-1}\left(\frac{k-i}{n-i}\right)^{2}\sum_{\begin{subarray}{c}\alpha\in\mathbb{N}^{n\times d}\\ |\alpha|=m\\ |\mathsf{RS}(\alpha)|=\ell\end{subarray}}\frac{1}{\left|[\alpha]\right|}$ (143)
$=1+\sum_{m=1}^{\infty}\rho^{2m}\sum_{\ell=0}^{k}\prod_{i=0}^{\ell-1}\left(\frac{k-i}{n-i}\right)^{2}\left|\left\{[\alpha]~{}\bigg{|}\begin{subarray}{c}\alpha\in\mathbb{N}^{n\times d}\\ |\alpha|=m\\ |\mathsf{RS}(\alpha)|=\ell\end{subarray}\right\}\right|.$ (144)
@@
We now turn to estimate the number of equivalence classes $\mathsf{[}\alpha]$, for $\alpha\in\mathbb{N}^{n\times d}$, $|\alpha|=m$ and $|\mathsf{RS}(\alpha)|=\ell$, for a fixed $\ell\leq k$. To that end, as in the one-dimensional case, we notice to a correspondence between equivalence classes and $d$-dimensional integer distribution functions with certain properties. Specifically, for $\alpha\in\mathbb{N}^{n\times d}$, consider the function $p_{\alpha}:\mathbb{N}^{d}\to\mathbb{N}$, such that $p_{\alpha}(v)$ is the number of rows of $\alpha$, which are equal to $v\in\mathbb{N}^{d}$. We also note that $|\alpha|=m$ and $|\mathsf{RS}(\alpha)|=\ell$ if and only if
@@
where $p_{1}$ is a partition of the integer $m$ to exactly $\ell$ parts, [^2] It is convenient to think about $p_{1}$ as a one-dimensional integer distribution function, such that $p_{1}(q)$ is the number of parts of size $q$ in a given partition. and,
[…]
$\leq d^{m}|\mathsf{Par}(m,\ell)|.$ (148)
@@
[…]
Proof of (147). We upper bound the number of the unordered lists described above by using a procedure that generates all possible lists, and then bounding the number of distinct lists generated by this procedure. We first go over all possible norms of vectors $v$ in that list. Since the summation of these weights is exactly $m$, we should go over all partitions of the integer $m$ to exactly $\ell$ elements. Having the norms of the vectors in the list fixed, we upper bound the number of ways to choose integer vectors with these exact norms. For a partition part with weight $q$, the number of vectors in $\mathbb{N}^{d}$ with weight $q$ is exactly the same as the number of ways to distribute $q$ elements into $d$ cells, which is known to be equal to $\binom{q+d-1}{d-1}$. For a fixed partition $p_{1}$, there are $p_{1}(q)$ parts with weight $q$, and therefore, there are at most $\binom{q+d-1}{d-1}^{p_{1}(q)}$ ways to generate $d$-dimensional integer vectors with such norm. Taking into account all of the distinct parts of the partition $p_{1}$, we upper bound the number ways to generate appropriate $d$-dimensional integer vectors by,
@@
Proof of (148). For the second upper bound, we note that the factor $\binom{q+d-1}{d-1}$, which is the number of ways to distribute $q$ elements into $d$ cells, can be upper bounded by $d^{q}$. If each one of the $q$ elements is a associated with a distinct item, they have $d$ optional cells each, and therefore, together, the $q$ elements have $d^{q}$ optional distributions along the $d$ cells. Hence,
@@
Proof of (146). For the lower bound, we note that for a fixed $q$, the expression in (149) may over-count the number of possible ways to choose $p_{1}(q)$ unordered $d$-dimensional vectors whose norms are $q$, because different orderings of such vectors are counted multiple times. Thus, we lower bound the number of possible ways to choose $p_{1}(q)$ unordered $d$-dimensional vectors with norms equal to $q$, by dividing by $p_{1}(q)!$, the maximal number of multiple counts for each set of $p_{1}(q)$ vectors. The rest of the analysis is carried out similarly to the upper bound. ∎
@@
Proofs of Theorems 56.
We continue from (144), and use Lemma 4 to obtain,
$\mathbb{E}_{{\cal H}_{0}}\left[{\cal L}(\mathsf{X},\mathsf{Y})^{2}\right]$ $=1+\sum_{m=1}^{\infty}\rho^{2m}\sum_{\ell=0}^{k}\prod_{i=0}^{\ell-1}\left(\frac{k-i}{n-i}\right)^{2}\left|\left\{[\alpha]~{}\bigg{|}\begin{subarray}{c}\alpha\in\mathbb{N}^{n\times d}\\ |\alpha|=m\\ |\mathsf{RS}(\alpha)|=\ell\end{subarray}\right\}\right|$ (155)
$=1+\sum_{m=1}^{\infty}\rho^{2m}\sum_{\ell=0}^{k}\prod_{i=0}^{\ell-1}\left(\frac{k-i}{n-i}\right)^{2}\left|\left\{[\alpha]~{}\bigg{|}\begin{subarray}{c}\alpha\in\mathbb{N}^{n\times d}\\ |\alpha|=m\\ |\mathsf{RS}(\alpha)|=\ell\end{subarray}\right\}\right|$ (156)
$=1+\sum_{m=1}^{\infty}\rho^{2m}\sum_{\ell=0}^{k}\prod_{i=0}^{\ell-1}\left(\frac{k-i}{n-i}\right)^{2}d^{m}\left|\mathsf{Par}(m,\ell)\right|$ (157)
$\overset{(a)}{\leq}1+\sum_{m=1}^{\infty}\left(d\rho^{2}\right)^{m}\sum_{\ell=0}^{k}\left(\frac{k}{n}\right)^{2\ell}\left|\mathsf{Par}(m,\ell)\right|$ (158)
$\overset{(b)}{\leq}1+\sum_{m=1}^{\infty}\left(d\rho^{2}\right)^{m}\sum_{\ell=0}^{k}\left(\frac{k}{n}\right)^{2}\left|\mathsf{Par}(m,\ell)\right|$ (159)
$=1+\left(\frac{k}{n}\right)^{2}\sum_{m=1}^{\infty}\left(d\rho^{2}\right)^{m}\sum_{\ell=0}^{k}\left|\mathsf{Par}(m,\ell)\right|$ (160)
$=1+\left(\frac{k}{n}\right)^{2}\sum_{m=1}^{\infty}\left(d\rho^{2}\right)^{m}\left|\mathsf{Par}(m,\leq_{k})\right|,$ (161)
where $(a)$ follows since the function $\frac{k-i}{n-i}$ is monotonically decreasing as a function of $i$, as long as $n>k$, and $(b)$ follows since $|\mathsf{Par(m,0)}|=0$, for any positive integer, and $(k/n)^{2\ell}$ is decreasing with $\ell$.
Now, recall the well-known formula for the generating function of integer partitions to at most $k$ elements (see, for example, [29, Chapter 1.3]),
@@
where the infinite sum converges when $|z|<1$. Thus, for any $\rho^{2}<\frac{1}{d}$, the infinite sum in (161) converges, and we have,
$\mathbb{E}_{{\cal H}_{0}}[{\cal L}(\mathsf{X},\mathsf{Y})^{2}]$ $\leq 1+\left(\frac{k}{n}\right)^{2}\sum_{m=1}^{\infty}|\mathsf{Par}(m,\leq_{k})|(d\rho^{2})^{m}$ (163)
$\leq 1-\left(\frac{k}{n}\right)^{2}+\left(\frac{k}{n}\right)^{2}\sum_{m=0}^{\infty}|\mathsf{Par}(m,\leq_{k})|(d\rho^{2})^{m}$ (164)
$=1+\left(\frac{k}{n}\right)^{2}\left(\prod_{i=1}^{k}\frac{1}{1-(d\rho^{2})^{i}}-1\right).$ (165)
The proof is concluded by repeating the same exact steps as in the proofs of Theorems 12, by using Cauchy-Schwartz inequality (32) for the impossibility of weak detection, and Lemma 4 for the impossibility of strong detection. ∎
Let us now compare our proof technique with the one in [4, 6, 7], where a straightforward analysis of $\mathbb{E}_{{\cal H}_{0}}[{\cal L}^{2}]$ is carried out. Repeating the same steps in [6], while generalizing the analysis for the partly correlated scenario, one can show that $\mathbb{E}_{{\cal H}_{0}}[{\cal L}(\mathsf{X},\mathsf{Y})^{2}]$ can be expressed using the following representation of random permutations cycle enumerators. The proof of the following result appear is relegated to Appendix .2.
@@
The second moment of the likelihood ratio for the detection problem defined in (LABEL:eqn:decproblem2) is given by
@@
In the fully correlated scenario where $k=n$, our generalized model reduces to the one studied in [6, 4]. In this case, $\mathsf{N}_{\ell}(\sigma,{\cal K})$ becomes $N_{\ell}(\sigma)$, which is the number of cycles of length $\ell$ of a random permutation. The analysis of (166) is already involved and requires a delicate case by case analysis for the different asymptotic regimes of $n,d$, and $\rho$, using a certain Poisson approximation of the random variables $(N_{\ell}(\sigma))_{\ell}$. Our method for calculating $\mathbb{E}_{{\cal H}_{0}}[{\cal L}(\mathsf{X},\mathsf{Y})^{2}]$, however, is able to recover the results in [6], but for all regimes at once.
In the case where $k<n$, it becomes unclear how to evaluate the expression in (166). Conditioning on the size of the intersection ${\cal K}\cap[k]$, which we denote here by $i$, it is possible to express (166) as a weighted average of elements of the form
@@
where $\mathsf{N}_{\ell}(\sigma,[i])$ is the size of the number of cycles of $\sigma$ intersecting the interval $[i]$ in exactly $\ell$ points. It seems likely that a Poisson approximation for the variables $(\mathsf{N}_{\ell}(\sigma,[i]))_{\ell,i}$ exists, and may be used in order to estimate $\mathbb{E}_{{\cal H}_{0}}[{\cal L}(\mathsf{X},\mathsf{Y})^{2}]$. However, proving such a Poisson approximation theorem is highly non-trivial, and has its own intellectual merit. We therefore leave it for future research.